Find the variance of the R.V. whose M.G.F is ((e-t)/12) * (2 + et + 6e3t + 3e6t)


1
Expert's answer
2021-09-29T07:23:33-0400
"Var(X)=\sigma^2=E[X^2]-(E[X])^2"

"=M''(0)-(M'(0))^2"

"M(t)=\dfrac{e^{-t}}{12}(2+e^t+6e^{3t}+3e^{6t})"

"=\dfrac{2e^{-t}+1+6e^{2t}+3e^{5t}}{12}"

"M'(t)=\dfrac{-2e^{-t}+12e^{2t}+15e^{5t}}{12}"

"M''(t)=\dfrac{2e^{-t}+24e^{2t}+75e^{5t}}{12}"

"M'(0)=\dfrac{-2+12+15}{12}=\dfrac{25}{12}"

"M''(0)=\dfrac{2+24+75}{12}=\dfrac{101}{12}"

"Var(X)=\sigma^2=\dfrac{101}{12}-(\dfrac{25}{12})^2"

"=\dfrac{1212-625}{144}=\dfrac{587}{144}"